IPP Progress Report for the week 2012.11.12 - 2012.11.17

(Up to IPP Progress Reports)

Eugene Magnier

  • NSF proposals finally finished. I made some interesting progress on parallax measurements in the DVO database in the process. I extracted the DVO/PS1 3pi parallaxes for the known L & T dwarfs and compared the results to the values determined from dedicated parallax programs. We agree quite well, with a chisq value (for the comparison) of ~1.7, telling us that our error model is reasonable. I put some plots on the ps1sc wiki: http://ps1sc.ifa.hawaii.edu/PS1wiki/index.php/DRAVG_telecon.

Serge Chastel

  • Czar on Tue and Wed
  • Cluster Storage:
    • Modified ~ipp/local/bin/cluster_status.py (see script and Cluster Storage Notes wikipage)
    • Added skycal to the list of stages
  • MOPS:
    • More data for test of DV3 format
    • Met with Larry and Peter
    • Set up google doc spreadsheet to keep track of MOPS various requests (reprocessing, ICD modification, bugs...)
  • Condor: analyzing a bug (or a Condor feature) I don't understand yet. Looks like setting a value for the Rank (which is used to tell condor to execute preferably on some targets) can prevent execution. I wonder if it's not a conifguration problem. Well... I'm still investigating.
  • ippdb04 (aka nebulous backup backup): at ATRC. Haydn/Steph/Rita are installing it.

Heather Flewelling

  • vacation

Mark Huber

  • holiday czar monday
  • sick day
  • continuing to trace out y magnitude offset and scatter problem in WS diffim reported by MOPS
  • MD04 prep for 2012/2013 season

Bill Sweeney

  • spent most of this week working on enhancements to cleanup that will recover more space and do proper logging.
  • Found a number of areas where small amounts of space have leaked through cleanup. For example there are some number of runs marked as clean that still have images on disk.
  • "arranged" to get 3 OSS exposures with seeing worse than the allowed 3 arcseconds to get through processing. MOPS had some simultaneous observations with another telescope that night and they wanted the images to complete processing.
  • extracted some information about all of the standard star observations from the database for John T.
  • Built some sample PSPS queries that show how to use the flags in the new schema for Ken. (64 bit numbers are a pain with sql server!)
  • looked into the cause of some out of range values in the SAS 10 Object table. Looks like DVO is not initializing some columns in some cases.
  • packaged some chip images for the outreach program.
  • debugged some postage stamp server problems. (Mostly caused by wave 1 node hangs)

Chris Waters

  • Finished diff analysis wiki page, and sent it to Larry and MOPS for comments.
  • Czaring duties and clearing out of old LAP runs that had become stuck for various reasons.
  • Debugging ppStackMedian. This now works, although there is one remaining bug to fix that causes the program to believe it has failed when it hasn't.